OpenStudy (anonymous):

Find the derivative of f(x) = negative 3 divided by x at x = -4.

OpenStudy (anonymous):

so 3/4?

OpenStudy (danjs):

nevermind , i am way off, and was doing the other way , integrate

OpenStudy (danjs):

[ 3 / x] ' = you can use the quotient derivative rule, or just rewrite the thing as a negative exponent and do product rule... [ 3*x^(-1) ] ' = -1*3*x^(-2)

OpenStudy (anonymous):

i am so confused now

OpenStudy (danjs):

then put the -2 exponent back -3 / x^2

OpenStudy (danjs):

oh the prob is -3, so the sign is + sorry f(x) = -3/x = -3*x^(-1) f ' (x) = 3*x^(-2) = 3/ x^2

OpenStudy (anonymous):

so 3/16?

OpenStudy (danjs):

the first kinda shortcut rule you learn , derivative of a variable to a power x^n \[\large \frac{ d }{ dx }x^n = n * x^{n-1}\]

OpenStudy (danjs):

yeah f ' (-4)

OpenStudy (danjs):

unless you have to show something else you learned,

OpenStudy (anonymous):

wait i thought the answer is 3/4

OpenStudy (danjs):

the derivative is f '(x) = 3/x^2 , evaluated at x=-4, f '(-4) = 3/16